OpenStudy (anonymous):

Choose the ordered pair (x,y) that is a solution of the equation. -8x + 5y = 96

OpenStudy (amoodarya):

where are choices ? anyway you can put x and find y if ok choice

OpenStudy (anonymous):

-3,11) (1,-19) (3,19) (-2,16)

OpenStudy (amoodarya):

for example put (1,-19) into the line -8x + 5y = 96 -8(1) +5(-19)=-8-95 ? =96 its not a correct choice now do like that

OpenStudy (anonymous):

thanks